OpenStudy (anonymous):

4x^2/(x^2-4) need domain and range... Im confused with this one

OpenStudy (anonymous):

I thought domain was (-INF,-2)U(-2,2)U(2,INF) and range (-INF,INF) but its not, why?

OpenStudy (anonymous):

it skips some stuff

OpenStudy (anonymous):

domain is easy to find, range is a bear

OpenStudy (anonymous):

set \[y=\frac{4x^2}{x^2-4}\] and solve the equation for \(x\) you need help doing that?

OpenStudy (anonymous):

another way would be to look at the graph it has a horizontal asymptote at \(y=4\) and vertical ones at \(x=-2\) and \(x=2\) in the interval \((-2,2)\) the curve goes from \(-\infty\) up to \(0\) then it skips up to the horizontal asymptote of \(y=4\) so the range is actually \[(-\infty, 0]\cup (4,\infty)\]

OpenStudy (anonymous):

I missed the bracket on zero in my original attempt.....what about the crazy domain

OpenStudy (anonymous):

the range is correct

OpenStudy (anonymous):

your domain is correct

OpenStudy (anonymous):

the range is what i wrote, not \((-\infty, \infty)\)